# Ensemble Kalman Filtering-Aided Variational Inference for Gaussian Process State-Space Models

### 23 Dec 2023 v2

@@ Abstract
Gaussian process state-space models (GPSSMs) are a flexible and principled approach for modeling dynamical systems. However, existing variational learning and inference methods for GPSSMs often necessitate optimizing a substantial number of variational distribution parameters, leading to inadequate performance and efficiency. To overcome this issue, we propose incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the variational inference framework to approximate the posterior distribution of latent states. This utilization of EnKF can effectively exploit the dependencies between latent states and GP dynamics, while eliminating the need for parameterizing the variational distribution, thereby significantly reducing the number of variational parameters. Moreover, we show that our proposed algorithm allows straightforward evaluation of an approximated evidence lower bound (ELBO) in variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging automatic differentiation tools, we hence can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed algorithm to an online setting and provide detailed algorithmic analyses and insights. Extensive evaluation on diverse real and synthetic datasets demonstrates the superiority of our EnKF-aided variational inference algorithms in terms of learning and inference performance compared to existing methods.
@@ I Introduction
State-space models (SSMs) describe the underlying dynamics of latent states through a transition function and an emission function [1]. As a versatile tool for modeling dynamical systems, SSM finds successful applications in diverse fields, including control engineering, signal processing, computer science, and economics [2, 3, 4]. In SSM, a key task involves inferring unobserved latent states from a sequence of noisy measurements. Established techniques, such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF), have been widely employed over the decades for latent state inference [2]. However, these state inference methods heavily rely on precise knowledge of the underlying system dynamics [5], posing tremendous challenges in complex and uncertain scenarios like model-based reinforcement learning [6, 7] and disease epidemic propagation [8]. Consequently, the underlying dynamics need to be learned from noisy observations, leading to the emergence of data-driven state-space models. One class of prominent data-driven state-space models is the Gaussian process state-space model (GPSSM) [9], which utilizes Gaussian processes (GPs) as core learning modules to capture the complex underlying system dynamics [10].
@@
Gaussian processes, serving as a Bayesian nonparametric model [11], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate model overfitting [12], rendering them more effective in scenarios with limited data [13]. Moreover, inheriting the Bayesian nature from GPs, GPSSMs offer interpretability and explicit uncertainty calibration for the system dynamics. As a result, GPSSMs exhibit increased transparency and caution in regions of the state space where data is scarce, which is crucial for safety-critical applications [14]. Due to these advantageous properties, GPSSMs have seen extensive use in various applications, such as human pose and motion learning [14], robotics and control learning [15, 16, 17], reinforcement learning [18], target tracking and navigation [19], and magnetic-field estimation [20].
@@
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the quality of state estimates affects model learning and vice versa, which introduces computational and statistical complexities. Second, the nonlinearity in GPSSMs renders the learning and inference analytically intractable [11]. Hence, the main goal in GPSSM is to accurately approximate the joint posterior over latent states and GP-modeled system dynamics. Multiple approaches have emerged towards achieving this objective in the last decade.
@@
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) is proposed in [9]. Subsequently, more advanced methods [21, 22, 23] leverage the reduced-rank GP approximation introduced in [24] to alleviate the substantial computational demands of GPs in [9]. However, the computational burden of the involved PMCMC remains notably significant, particularly when dealing with high-dimensional latent states. Consequently, there has been a shift towards variational inference methods [25, 26, 27, 28, 29, 30, 31, 32, 33], which utilize the sparse GP approximations with inducing points [34].
@@
Generally, the variational inference approaches can be classified into two categories: the mean-field (MF) class [25, 26, 27, 28], and the non-mean-field (NMF) class [29, 30, 31, 32, 33], based on whether the independence assumption is applied to the variational distribution concerning latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, is introduced in [25]. Subsequent MF-based endeavors [26, 27, 28] aim at reducing computational and model complexities in [25]. Yet, the MF class method neglects the dependencies between latent states and GP dynamics, potentially compromising learning accuracy and yielding overconfident state estimates [35].
@@
More recent methods concentrate on NMF variational approximations [29, 30, 31, 32, 33]. Specifically, in [29], the posterior distribution of underlying states conditioned on the GP dynamics is directly approximated by the corresponding prior distribution; while in [30] and [31], the posterior is approximated by parametric Markov-structured Gaussian variational distributions. In contrast to the heavy parameterization as in [30] and [31], the works in [32] and [33] opt for directly characterizing the posterior distribution through the model evidence. However, the Laplace approximation in [32] can oversimplify the approximated distribution over latent states (unimodal approximation) [11], and the stochastic gradient Hamiltonian Monte Carlo method [36] employed in [33] exhibits slow convergence ($50,000$ iterations as reported). Other recent works, such as [37, 38, 39, 40], have highlighted the advantageous impact of increased GPSSM’s flexibility on learning and inference performance.
Nevertheless, the existing methods face two significant issues when approximating the joint posterior distribution concerning latent states and GP dynamics. First, the majority of existing approaches [27, 28, 29, 30, 31, 33, 38, 39, 40] continue to struggle with computational challenges, manifested by the extensive optimization parameters and notably slow convergence of the optimization algorithms. Additionally, the existing variational methods are incompatible with online applications, posing a notable limitation in their applicability.
To address the two issues mentioned above, our proposed methods focus on two critical aspects in terms of the joint posterior distribution approximation: 1) whether the dependencies between GP dynamics and latent states are fully exploited during algorithm design and 2) whether the corresponding variational distributions are as less parameterized as possible. We show that by exploiting the dependencies between latent states and GP dynamics via EnKF, our methods can significantly alleviate the model and computational complexity and get readily extended to the online setting. The main contributions of this paper are summarized as follows.
@@
  • We highlight the significance of incorporating well-established model-based filtering techniques to enhance the efficiency of GPSSM variational algorithms. In contrast to existing methods utilizing numerous variational parameters to parameterize the variational distribution over latent states, our proposed novel EnKF-aided NMF variational inference algorithm harnesses the EnKF to exploit the dependencies between latent states and GP dynamics, and eliminate the need to parameterize the variational distribution, significantly reducing the number of variational parameters and accelerating the algorithm convergence.
@@
  • By leveraging EnKF, we also demonstrate the simplified computation of the approximated evidence lower bound (ELBO) in variational inference, by aggregating multiple terms with readily available closed-form analytical solutions. Leveraging the differentiability inherent in EnKF alongside automatic differentiation tools, we can efficiently optimize the ELBO and train the GPSSM.
@@
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to an online learning setting. This online extension introduces a more comprehensive objective function, making it superior to state-of-the-art online algorithms in learning and inference efficiency.
@@
  • We provide detailed analysis and fresh insights of the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the algorithms’ performance. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform existing state-of-the-art methods in terms of learning and inference performance.
@@
The remainder of this paper is organized as follows. Some preliminaries related to GPSSMs are provided in Section II. The problems addressed in this paper are summarized in Section III. Section IV elaborates the proposed EnKF-aided variational learning algorithm. Section V extends the proposed algorithm to the online setting. The experimental results are presented in Section VI, and Section VII concludes this paper. More supportive results, proofs, etc. are relegated to the Appendix and supplementary materials.
@@ II-A Gaussian Processes (GPs)
A GP defines a collection of random variables indexed by $\bm{x}\in{\cal X}$, such that any finite subset of these variables follows a joint Gaussian distribution [10]. Mathematically, a real scalar-valued GP $f(\bm{x})$ can be written as
@@
where $\mu({\bm{x}})$ is a mean function typically set to zero in practice, and $k(\bm{x},\bm{x}^{\prime})$ is the covariance function, also known as kernel function, which is interpretable and can provide insights about the nature of the underlying function [10]; and $\bm{\theta}_{gp}$ is a set of hyperparameters that needs to be tuned for model selection.
Let us consider a general regression model,
$y=f(\bm{x})+{e},\quad{e}\sim{\cal N}(0,\sigma_{e}^{2}),\quad y\in\mathbb{R}.$ (2)
By placing a GP prior over the function $f(\cdot):{\cal X}\mapsto\mathbb{R}$, we get the salient Gaussian process regression (GPR) model. The task in GPR model is to infer the mapping function $f(\cdot)$ using an observed dataset $\mathcal{D}\triangleq\{\bm{x}_{i},y_{i}\}_{i=1}^{n}\triangleq\{\bm{X},\bm{y}\}$ consisting of $n$ samples, or alternatively input-output pairs. Conditioning on the observed data, the posterior distribution of the mapping function, $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$, at any test input $\bm{x}_{*}\in{\cal X}$, is Gaussian, fully characterized by the posterior mean $\xi$ and the posterior variance $\Xi$. Concretely,
@@
where $\bm{K}_{\bm{X},\bm{X}}$ denotes the covariance matrix evaluated on the training input $\bm{X}$, and each entry is $[\bm{K}_{\bm{X},\bm{X}}]_{i,j}=k({\bm{x}}_{i},\bm{x}_{j})$; $\bm{K}_{\bm{x}_{*},\bm{X}}$ denotes the cross covariance matrix between the test input $\bm{x}_{*}$ and the training input $\bm{X}$; the zero-mean GP prior is assumed here and will be used in the rest of this paper if there is no further specification. Note that the posterior distribution $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$ gives not only a point estimate, i.e., the posterior mean, but also an uncertainty region of such estimate quantified by the posterior variance. It should also be noted that here we denote the variables in the GPR model using mathematical mode italics, such as $\bm{x}_{i}$ and $y_{i}$; these variables should not be confused with the latent state ${\mathbf{x}}_{t}$ and observation ${\mathbf{y}}_{t}$ in SSM (cf. Eq. (II-B)).
@@ II-B Gaussian Process State-Space Models (GPSSMs)
The GPSSM incorporates a GP prior to model the transition function $f(\cdot)$ in Eq. (II-B). Specifically, Fig. 1 presents the graphical model of GPSSM, while the following equations express its formulation:
@@
where $\mathbf{f}_{t}$ represents the GP transition function value evaluated at the previous state $\mathbf{x}_{t-1}$. Note that when dealing with a multidimensional state space ($d_{x}\!>\!1$), the transition function $f(\cdot):\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$ is represented using a multi-output GP. In this context, the $d_{x}$ functions are typically modeled with $d_{x}$ mutually independent GPs [37]. Furthermore, it is worth noting that the GPSSM illustrated in Fig. 1 can be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state with $[{\mathbf{x}}_{t},\bm{c}_{t}]\in\mathbb{R}^{d_{x}+d_{c}}$. For the sake of brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.
@@
where $p({\mathbf{f}}_{1:T})\!\!=\!\!p(f({\mathbf{x}}_{0:T-1}))\!\!=\!\!\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{f}}_{1:t-1},{\mathbf{x}}_{0:t-1})$ corresponds to a finite dimensional ($T$-dimensional) GP distribution [26], and we define the short-hand notations ${\vec{\mathbf{y}}}\triangleq{\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, ${\vec{\mathbf{f}}}\triangleq{\mathbf{f}}_{1:T}=\{{\mathbf{f}}_{t}\}_{t=1}^{T}$, and ${\vec{\mathbf{x}}}\triangleq{\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$. The model parameters $\bm{\theta}$ includes the noise covariance and GP hyper-parameters, i.e., $\bm{\theta}\!\!=\!\![\mathbf{Q},\mathbf{R},\bm{\theta}_{gp}]$. The challenging task in GPSSM is to learn ${{\bm{\theta}}}$, and simultaneously infer the latent states of interest, which usually involves the marginal distribution $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. However, due to the nonlinearity of GP, a closed-form analytical solution for $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is unavailable. Hence, it becomes necessary to utilize approximation methods.

### 08 Jan 2024 v3

# Ensemble Kalman Filtering Meets Gaussian Process SSM for Non-Mean-Field and Online Inference

@@ Abstract
Gaussian process state-space models (GPSSMs) are a versatile and principled family of nonlinear dynamical system models. However, existing variational learning and inference methods for GPSSMs often necessitate optimizing a substantial number of variational parameters, leading to inadequate performance and efficiency. To overcome this issue, we propose incorporating the ensemble Kalman filter (EnKF), a well-established model-based filtering technique, into the variational inference framework to approximate the posterior distribution of latent states. This utilization of EnKF can effectively exploit the dependencies between latent states and GP dynamics, while eliminating the need for parameterizing the variational distribution, thereby significantly reducing the number of variational parameters. Moreover, we show that our proposed algorithm allows straightforward evaluation of an approximated evidence lower bound (ELBO) in variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging automatic differentiation tools, we hence can maximize the ELBO and train the GPSSM efficiently. We also extend the proposed algorithm to accommodate an online setting and provide detailed algorithmic analyses and insights. Extensive evaluation on diverse real and synthetic datasets demonstrates the superiority of our EnKF-aided variational inference algorithms in terms of learning and inference performance compared to existing methods.
@@ I Introduction
State-space models (SSMs) describe the underlying dynamics of latent states through a transition function and an emission function [1]. As a versatile tool for modeling dynamical systems, SSM finds successful applications in diverse fields, including control engineering, signal processing, computer science, and economics [2, 3, 4]. In SSM, a key task is to infer unobserved latent states from a sequence of noisy measurements. Established techniques, such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF), have been widely employed over the decades for latent state inference [2]. However, these classic state inference methods heavily rely on precise knowledge of the underlying system dynamics [5], posing tremendous challenges in complex and uncertain scenarios like model-based reinforcement learning [6] and disease epidemic propagation [7]. As a novel alternative, the underlying complex dynamics can be learned from noisy observations, leading to the emergence of data-driven SSMs. One class of prominent data-driven SSMs is the Gaussian process state-space model (GPSSM) [8], which utilizes Gaussian processes (GPs) [9] as the core learning modules to capture the complex underlying system dynamics.
@@
Gaussian processes, serving as the most prominent Bayesian nonparametric model [10], provide the flexibility to model general nonlinear system dynamics without enforcing an explicit parametric structure. With the inherent regularization imposed by the GP prior, GPSSMs are able to mitigate model overfitting and generalization issues [11], rendering them to be more effective in scenarios with limited data samples [12]. Moreover, owing to its Bayesian nature, GPSSMs maintain good interpretability and explicit uncertainty calibration for analyzing system dynamics. These superior properties have led to the extensive usage of GPSSMs in various live applications, such as human pose and motion learning [13], robotics and control learning [14], reinforcement learning [15], target tracking and navigation [16], and magnetic-field sensing [17].
@@
Despite the popularity of GPSSMs, simultaneously learning the model and estimating the latent states in GPSSMs remains highly challenging, primarily due to the following two factors. First, the quality of the state estimates affects the model learning and vice versa, leading to heightened computational and statistical complexities. Second, the nonlinearity in GPSSMs prohibits tractable learning and inference process [10]. Hence, the main task in GPSSM is to accurately approximate the joint posterior over the latent states and the system dynamics represented by GPs. Multiple approaches have emerged towards achieving this objective in the last decade.
@@
The seminal work utilizing particle Markov chain Monte Carlo (PMCMC) was proposed in [8]. Subsequently, more advanced methods [18, 19, 20, 21] leveraged the reduced-rank GP approximation introduced in [22] to alleviate the substantial computational demands of GPs in [8]. However, the computational burden of the involved PMCMC remains unaffordable, particularly when dealing with high-dimensional latent states. Consequently, there has been a paradigm shift towards variational inference methods [23, 24, 25, 26, 27, 28, 29, 30, 31], which adopted the classic sparse GP approximations with inducing points [32].
@@
Generally, the variational inference approaches can be classified into two main categories: the mean-field (MF) class [23, 24, 25, 26] versus the non-mean-field (NMF) class [27, 28, 29, 30, 31], depending on whether the statistical independence assumption has to be applied to the variational distribution of the latent states and GP dynamics. The first variational algorithm, incorporating the MF assumption and integrating PF for latent state inference, was introduced in [23]. Subsequent MF endeavors [24, 25, 26] aimed at reducing computational and model complexities in [23]. Yet, the main idea of the MF class is to neglect the inherent dependencies between the latent states and the GP dynamics, potentially impairing the overall learning accuracy and yielding overconfident state estimates [33].
@@
Recent methods [27, 28, 29, 30, 31], therefore, concentrate on the more complicated NMF approximation to address the inaccuracies arising from the MF assumption, yet they continue to struggle with computational challenges. Specifically, in [27], the posterior distribution of underlying states conditioned on the GP dynamics was simply approximated by the subjectively selected prior distribution, while in [28] and [29], the posterior was approximated by some parametric Markov-structured Gaussian variational distributions. However, optimizing the extensive variational parameters in the variational distributions is computationally demanding and inefficient. In contrast to the heavy parameterization as conducted in [27, 28, 29], the works in [30] and [31] opted for directly characterizing the posterior distribution through the model evidence. Nonetheless, the Laplace approximation, employed in [30], can yield an oversimplified unimodal posterior distribution over the latent states [10], while the stochastic gradient Hamiltonian Monte Carlo method [34], employed in [31], may exhibit very slow convergence ($50,000$ iterations as reported). Furthermore, all these existing variational methods are incompatible with online applications, posing a notable limitation in their applicability.
To tackle the computational challenges mentioned above, our proposed methods incorporate the model-based filtering technique, EnKF, to approximate the joint posterior distribution. By leveraging the EnKF, our method can fully exploit the dependencies between the latent states and the GP dynamics, while reducing the parameterization of associated variational distributions, thereby alleviating model and computational complexities and accelerating algorithm convergence. This streamlined parameterization also helps the easy accommodation of our method in an online setting. The main contributions of this paper are summarized as follows.
@@
  • We corroborate the significance of incorporating well-established model-based filtering techniques to enhance the efficiency of GPSSM variational algorithms. In contrast to existing methods utilizing numerous variational parameters to parameterize the variational distribution over latent states, our proposed novel algorithm harnesses the EnKF to accurately capture the dependencies between the latent states and the GP dynamics, and eliminate the need to parameterize the variational distribution, significantly reducing the number of variational parameters and accelerating the algorithm convergence.
@@
  • By leveraging EnKF, we also demonstrate a simplified evidence lower bound (ELBO) for variational inference via simply summating multiple terms with readily available closed-form solutions. Leveraging the differentiable nature of the classic EnKF alongside the off-shelf automatic differentiation tools, we can optimize the ELBO and train the GPSSM efficiently.
@@
  • Without explicitly parameterizing the variational distribution over latent states, the proposed EnKF-aided algorithm gets readily extended to accommodate online applications. This online method relies on a more principled objective by ensuring observation fitting accuracy while incorporating model regularizations to mitigate model overfitting, rendering it superior to state-of-the-art online algorithms in learning and inference efficiency.
@@
  • We provide detailed analysis and fresh insights for the proposed algorithms and conduct extensive experiments on various real and synthetic datasets to evaluate the performance of these algorithms from various aspects. The results demonstrate that the proposed EnKF-aided variational learning methods consistently outperform the existing state-of-the-art methods, especially in terms of learning and inference.
@@
The remainder of this paper is organized as follows. Some preliminaries related to GPSSMs are provided in Section II. The computational challenges of the existing methods are summarized in Section III. Section IV elaborates the proposed EnKF-aided variational learning algorithm. Section V extends the proposed algorithm to accommodate online applications. The experimental results are presented in Section VI, and Section VII concludes this paper. More supportive results, proofs, etc. are relegated to the Appendix and supplementary materials.
@@ II-A Gaussian Processes (GPs)
A GP defines a collection of random variables indexed by $\bm{x}\in\mathcal{X}$, such that any finite subset of these variables follows a joint Gaussian distribution [9]. Mathematically, a real scalar-valued GP $f(\bm{x})$ can be represented as
@@
where $\mu({\bm{x}})$ is a mean function typically set to zero in practice, and $k(\bm{x},\bm{x}^{\prime})$ is the kernel function that provides insights about the nature of the underlying function [9]; and $\bm{\theta}_{gp}$ is a set of hyperparameters that needs to be tuned for model selection. By placing a GP prior over the function $f(\cdot):{\cal X}\mapsto\mathbb{R}$ in a general regression model,
$y=f(\bm{x})+{e},\quad{e}\sim{\cal N}(0,\sigma_{e}^{2}),\quad y\in\mathbb{R},$ (2)
we get the salient Gaussian process regression (GPR) model.
Given an observed dataset, $\mathcal{D}\!\triangleq\!\{\bm{x}_{i},y_{i}\}_{i=1}^{n}\!\triangleq\!\{\bm{X},\bm{y}\}$ consisting of $n$ input-output pairs, the posterior distribution of the mapping function, $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$, at any test input $\bm{x}_{*}\!\in\!{\cal X}$, is Gaussian [9], fully characterized by the posterior mean $\xi$ and the posterior variance $\Xi$. Concretely,
@@
where $\bm{K}_{\bm{X},\bm{X}}$ denotes the covariance matrix evaluated on the training input $\bm{X}$, and each entry is $[\bm{K}_{\bm{X},\bm{X}}]_{i,j}=k({\bm{x}}_{i},\bm{x}_{j})$; $\bm{K}_{\bm{x}_{*},\bm{X}}$ denotes the cross covariance matrix between the test input $\bm{x}_{*}$ and the training input $\bm{X}$; the zero-mean GP prior is assumed here and will be used in the rest of this paper if there is no further specification. Note that the posterior distribution $p(f(\bm{x}_{*})|\bm{x}_{*},{\cal D})$ gives not only a point estimate, i.e., the posterior mean, but also an uncertainty region of such estimate quantified by the posterior variance. It is also noteworthy that here we denote the variables in the GPR using mathematical mode italics, such as $\bm{x}_{i}$ and $y_{i}$; these variables should not be confused with the latent state ${\mathbf{x}}_{t}$ and observation ${\mathbf{y}}_{t}$ in SSM (cf. Eq. (II-B)).
@@ II-B Gaussian Process State-Space Models (GPSSMs)
The GPSSM incorporates a GP prior to model the transition function $f(\cdot)$ in Eq. (II-B). Specifically, Fig. 1 presents the graphical model of GPSSM, while the following equations express its mathematical representation:
@@
where $\mathbf{f}_{t}$ represents the GP transition function value evaluated at the previous state $\mathbf{x}_{t-1}$. Note that when dealing with a multidimensional state space ($d_{x}\!>\!1$), the transition function $f(\cdot):\mathbb{R}^{d_{x}}\!\mapsto\!\mathbb{R}^{d_{x}}$ is represented using a multi-output GP. In this context, the $d_{x}$ functions are typically modeled with $d_{x}$ mutually independent GPs [36]. Furthermore, it is noteworthy that the GPSSM illustrated in Fig. 1 can be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state with $[{\mathbf{x}}_{t},\bm{c}_{t}]\in\mathbb{R}^{d_{x}+d_{c}}$. For the sake of brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.
@@
where $p({\mathbf{f}}_{1:T})\!\!=\!\!p(f({\mathbf{x}}_{0:T-1}))\!\!=\!\!\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{f}}_{1:t-1},{\mathbf{x}}_{0:t-1})$ corresponds to a finite dimensional ($T$-dimensional) GP distribution [24], and we define the short-hand notations ${\vec{\mathbf{y}}}\triangleq{\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, ${\vec{\mathbf{f}}}\triangleq{\mathbf{f}}_{1:T}=\{{\mathbf{f}}_{t}\}_{t=1}^{T}$, and ${\vec{\mathbf{x}}}\triangleq{\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$. The model parameters $\bm{\theta}$ includes the noise covariance and GP hyper-parameters, i.e., $\bm{\theta}\!\!=\!\![\mathbf{Q},\mathbf{R},\bm{\theta}_{gp}]$. The challenging task in GPSSM is to learn ${{\bm{\theta}}}$, and simultaneously infer the latent states of interest, which involves the marginal distribution $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. However, due to the nonlinearity of GP, a closed-form solution for $p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ is unavailable. Hence, it becomes necessary to utilize approximation methods.